OpenStudy (anonymous):

Evaluate 27 × a^-2 for a = 3.

OpenStudy (anonymous):

1/3 1/9 9 3

OpenStudy (anonymous):

Plug in 3 for a.

OpenStudy (anonymous):

Once you plug that in start solving.

OpenStudy (anonymous):

@Future_UsArmy_MP

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

What does 27 * 3^-2 equal to?

OpenStudy (anonymous):

Simple math, we're talking about here.

OpenStudy (anonymous):

and i want a simple explanation, thats what Open study is looking for

OpenStudy (anonymous):

You could solve this I gave you the equation. All you have to do is plug it in a caluclator.

OpenStudy (igreen):

\[27 \times a^{-2}\] \[27 \times 3^{-2}\] \[27 \times 0.11111111111111\] Can you multiply that? @Future_UsArmy_MP

OpenStudy (anonymous):

Thank you @iGreen. Great explanation.

OpenStudy (igreen):

Np